OpenStudy (anonymous):

cot (x-(pi/2))= -tan x Verify the identity.

OpenStudy (anonymous):

@nelsonjedi

OpenStudy (aum):

I suppose you can write it as cosine / sine, expand using the relevant identities and simplify.

OpenStudy (anonymous):

umm...

OpenStudy (aum):

cos(A-B) = cos(A)cos(B) + sin(A)sin(B) sin(A-B) = sin(A)cos(B) - cos(A)sin(B)

OpenStudy (anonymous):

thats is what i have for these problems

OpenStudy (aum):

That link won't work for people who don't have an account there.

OpenStudy (anonymous):

Reciprocal Identities sin u equals 1 over csc u cos u equals 1 over sec u tan u equals 1 over cot u csc u equals 1 over sin u sec u equals 1 over cos u cot u equals 1 over tan u Quotient Identities tan u equals sin u over cos u cotangent u equals cosine u over sine u Pythagorean Identities sin2u + cos2u = 1 1 + tan2u = sec2u 1 + cot2u = csc2u Cofunction Identities sin (pi over 2) equals cos u cos (pi over 2 minus u) equals sin u tan (pi over 2 minu u) equals cot u cot (pi over 2 minus u) equals tan u sec (pi over 2 minus u) equals csc u csc (pi over 2 minus u) equals sec u Even/Odd Identities sin(-u) = -sinu sec(-u) = secu tan(-u) = -tanu csc(-u) = -cscu cos(-u) = cosu cot(-u) = -cotu

OpenStudy (anonymous):

@aum

OpenStudy (anonymous):

but it has to equal -tanx @zzr0ck3r

OpenStudy (zzr0ck3r):

\(\frac{cos(x-\frac{\pi}{2})}{sin(x-\frac{\pi}{2})}=\frac{sin(x)}{-cos(x)}=-tan(x)\)

OpenStudy (zzr0ck3r):

fixed

OpenStudy (aum):

Okay, they don't want you to PROVE the identity but just verify it using the identities provided in the list above. One of the Cofunction identity states: cot (pi over 2 minus u) equals tan u That is, cot(pi/2 - u) = tan(u) or cot(pi/2 - x) = tan(x) Here, on the LHS. you have cot (x-(pi/2)) = cot ( - (pi/2-x) ) = -cot(pi/2 -x ) = -tan(x)

OpenStudy (zzr0ck3r):

@ninjasandtigers do you understand?

OpenStudy (anonymous):

yeah so we are basically just flipping the equation around a bit? @aum @zzr0ck3r

OpenStudy (zzr0ck3r):

its just the fact that \(\cos(x-\frac{pi}{2})=\sin(x)\) and \(\sin(x-\frac{\pi}{2})=-\cos(x)\)

OpenStudy (anonymous):

ok